The most recent financial statements for Schenkel Co are shown here Income Statement Balance Sheet Sales i Costs $17,600 Current assets $10,700 Debt $15,200 12.000 Fixed assets 25,750Equity21250 Taxable income 5600 Total $36 450Total 36,450 ーーーゴaxes (40%):t if 2.240 Net income 3360 Aasest fhiudentd pavoit fatiol Ma budtsna simanhcino is ausiblile not Tthe company maintains k eanstant to percent dividend payout ratio. No external financing is possible What tis the internal growth rate? Do not round intermediate calculations. Enter your answer as a ercent rounded 2 decimal places e g. 32 16:) Internal growth rate
Answer #1

ROA=Net income/Total assets

=(3360/36450)=0.092181069

Retention ratio=1-payout ratio

=(1-0.4)=0.6

Internal growth rate=(ROA*Retention ratio)/[1-(ROA*Retention ratio)]

=(0.092181069*0.6)/[1-(0.092181069*0.6)]

=5.85%(Approx).